Rigo:

In order to make an A on her project, Sarah needs at least 160 points. She has already turned in part of her work and has been given 125 point so far. Write and solve an inequality that models this situation and find out how many points Sarah needs on the last part of her project to get an A. A) at least 35 points B) at least 45 points C) less than 35 points D) less than 45 points

Razor:

Well, this seems simple, taking 160 and subtracting 125 from it should give you the answer since it's multiple choice. Unless you really need to write an equation

Rigo:

an equation4

Razor:

\[125 + P \ge 160\]

Rigo:

thx

Razor:

\[P \ge 160 - 125 \] Which then gives you \[P \ge 35\]